OpenStudy (anonymous):

How do I find the normalization constant? I've used variable elimination to calculate a conditional probability, and the textbook says the result is: a(.0003699, .001717)= (.1772, .8228) and a is the normalization constant. So how do I find a?

OpenStudy (sandra):

ah right hmmm. well the normalization constant is the number you multiply the conditional probabilities by so that the sum of the probabilities is 1

OpenStudy (sandra):

as you can see, .1772 + .8228 = 1

OpenStudy (sandra):

so I'm pretty hazy on probability, but maybe a*.0003699 + a*.001717 = 1

OpenStudy (sandra):

0.0020869 * a = 1

OpenStudy (sandra):

a = 1/0.0020869

OpenStudy (anonymous):

omg

OpenStudy (anonymous):

Lol thank you haha I'm so tired ... wtf I'm stupid

OpenStudy (anonymous):

I've been confusing a = a probability variable

OpenStudy (sandra):

heh np. hope that's right lol!

OpenStudy (anonymous):

Yes it is correct. Thanks very much.